A disk is mounted on a slender rod as shown below
Answer #1

Let friction be f
then f*r = Ia
where r is radius of disc, I is moment of inertia and a is angular acceleration of disc
also, f*r = I'a'
where I' is moment of inertia of rod and a' is its angular acceleration

now, -1.75 = -2 + a'*t = -2 + fr*t/I'
and -1.75 = w + at = w + fr*t/I
comparing both
0.252*I' = [-1.75 - w]I
w = -0.252I'/I - 1.75 = -0.252*2*(20*4^2)/3*8*0.5^2 - 1.75 = -28.63 rad/s